Solve the equation by the method of your choice. Simplify solutions, if possible.
(3x + 5) + = 6

Answer by Anglophile(11)

(3X+5) = 6
3X = 1
x = 1/3
Check:
3(1/3) + 5=6
1 + 5 = 6
6 = 6